Two tiny conducting spheres are identical and carry charges of -24.1C and +61.2C. They are separated by a distance of 1.89 cm. (a) What is the magnitude of the force that each sphere experiences? (b) The spheres are brought into contact and then separated to a distance of 1.89 cm. Determine the magnitude of the force that each sphere now experiences.
Answer #1

a)

As you Know force between two charge

F=Kq1q2/r^2 ...1

here K=9×10^9 Nm^2/C^2 ,q1= -24.1 C ,q2=+61.2 C ,r=1.89 cm

r=1.89×10^(-2) m

So Magnitude of the force

F=9×10^9×24.1×61.2 / (1.89×10^(-2))^2

F=3716.1×10^13

F=3.716×10^16 N. Answer

b)

when sphere brought into contact and then separated so charge on the each sphere

. q1=q2=(-24.1+61.2)÷2

. q1=q2=18.55 C

now Magnitude of the force

F=9×10^9×18.55×18.55/(1.89×10^(-2))^2

F=866.98×10^13

F=8.67×10^15 N Answer